If X is a binomial random variable counting the number of successes in n = 5 Bernoulli trials, each with probability of success p = .2, find Pr[X = 2], correct to 4 decimal places.

A. .4000

B. .2048

C. .2000

D. .1024

E. .0512
Answer #1

Given that ,

p = 0.2

1 - p = 1 - 0.2 = 0.8

(a)n = 5

Using binomial probability formula ,

P(X = x) = (n C x) * px * (1 - p)n - x

P(X = 2) = (5 C 2) * (0.2)2 * (0.80)3

=0.2048

probability=0.2048